OpenStudy (jackfrostt3482):

Please help asap I will give medal and fan! What is 1/3 x pi ???

OpenStudy (jackfrostt3482):

@jabez177

OpenStudy (3mar):

\[\huge(\frac{ 1 }{ 3 })×\pi\] Is that what you mean?

OpenStudy (jackfrostt3482):

yes

OpenStudy (3mar):

You may assume that \[\huge \pi=\frac{ 22 }{ 7 }\] so the expression would be: \[\huge (\frac{ 1 }{ 3 })(\frac{ 22 }{ 7 })=\frac{ 22 }{ 21 }\] I hope I helped you.

OpenStudy (jackfrostt3482):

Thank you!

OpenStudy (3mar):

You are welcome! Any Help... Any Time... Did you get what you are looking for?

OpenStudy (3mar):

Any more questions?

OpenStudy (jackfrostt3482):

yes, hold on

OpenStudy (3mar):

If you don't mind can we start in new post, please?

OpenStudy (3mar):

I have less than 10 min to go!

OpenStudy (jackfrostt3482):

how do i put that into v =